# Halfway to the Peak: The JWST MIRI 5.6 micron number counts and source population

### v1

@@ 2.5 Photometric zero-point
Pre-launch estimates of the MIRI imager’s performance already predicted that its point spread function (PSF) has a ‘cruciform’ artifact that draws luminosity from the core into the wings of the Gaussian. In-flight measurements have shown that this artifact is prominent at shorter wavelengths like ours (Rigby et al., 2023), but is negligible at 10 m (Gáspár et al., 2021). The extent of this effect was only noted once in-flight assessments had been made. Since our data were obtained very early after the start of science operations, the FITS header keyword PHOTUJA2 provided for converting from the native units to physical photometry units was incorrect. The fiducial keyword translated to a photometric zero-point of 26.121 mag which resulted in us seeing a significant systematic offset in the photometry measured by MIRI F560W as compared to the Spitzer IRAC Ch3 data available for our fields.
This issue was also noted by the CEERS team and led them to calculate a photometric zero-point for the 5.6 m images of 25.701 in AB magnitudes (Papovich et al., 2023). However, the CEERS images have a pixel scale of 0.09 (Papovich et al., 2023) whereas ours is 0.11. We re-scale the CEERS calculated zero-point as in . This is the value we adopt in this paper which leads to much better agreement between the MIRI 5.6 m and IRAC channel 3 photometry for cross-matched sources (see Section 3.3).
@@ 3.1 SExtractor photometry
We used Source Extractor (Bertin, E. & Arnouts, S., 1996) version 2.28.0 for source detection and photometry. Table 2 lists our SExtractor parameter settings. The SEEINGFWHM keyword adopted here corresponds to the FWHM of the in-flight measured MIRI 5.6 m PSF (see section 4.1 for details). As our fiducial flux measurements, we adopt values produced with Kron aperture photometry, labeled by the FLUX_AUTO keyword by SExtractor. The photometric uncertainties are natively provided by SE and take into account the variance of the pixel values inside the aperture, the background value, and the effective detector gain.
@@ 3.2 Masking
We further clean the raw SExtractor catalogs by masking out bright stars as well as any particularly noisy parts of the images. To do so we created masks for each image. These masks include two components. One excludes obviously diffracted sources. The other avoids all edges (assuming an edge width of 5 pixels), where edges include the blocked regions of the Lyot coronograph. We also manually remove the noisy residual stripes near the bottom of the images (see right-hand panel in Figure 2). Figure 3 shows as an example the mask for the FLS1 field.
We also apply a quality cut on our catalogs, only keeping sources with a signal-to-noise ratio greater than 5. The combination of the masking and the quality cut reduces the raw SExtractor output of 4,089 sources to a total of 2,854 “reliable” sources (“reliable” meaning falling outside our defined masks and having SNR) among all our eight fields. Per field, we detect the most in field FLS3 (409 sources) and the least in field FLS5 (310 sources).
@@ 3.3 Astrometric & photometric verification
In order to perform astrometric and photometric verifications, we used the SExtractor detected sources in our two COSMOS fields and cross-matched them with the COSMOS2020 CLASSIC catalog (Weaver et al., 2022a). The COSMOS2020 source detection is based on a weighted combined image (Weaver et al., 2022a). We used a 1′′ matching radius and found a total of 582 COSMOS2020 cross-matches, 175 of which also have IRAC channel 3 detections. Overall, our MIRI data increase the fraction of COSMOS2020 with 5.6/5.8 m data by roughly 3 times. Note that our MIRI fields are within the deep part of the COSMOS2020 catalog and not near the edges and the COSMOS2020 positions rely on Gaia astrometric solutions (Weaver et al., 2022a). Figure 4 shows the comparison between the MIRI and COSMOS2020 source positions of the cross-matched sources. The offsets we find are small and well within the MIRI PSF at this wavelength. Therefore we do not explicitly apply astrometric corrections.
Figure 5 shows the comparison between the MIRI 5.6 m and the available IRAC channel 3 fluxes. As described in section 3.1, we use Kron aperture photometry for the 5.6 m fluxes. For the IRAC fluxes, we used the aperture corrected 2′′ aperture photometry from Weaver et al. (2022b). We overlay the 1:1 line. Overall, we find good agreement, although we have a median offset of 0.22 magnitudes.
@@ 3.4 Estimating fraction of fake sources
The cross-match with COSMOS2020 also gives us an upper limit on the fraction of the SExtractor detected sources that might be fake. Such sources are expected due to some of the residual artifacts in the images, that may have been missed in the masks described in Section 3.2. Overall we found COSMOS2020 counterparts for 582 of the 690 MIRI sources in our two COSMOS fields (84 %). The 108 unmatched sources (), represent an upper limit on the fake source fraction since it is expected that we have some red MIRI sources that are not detected in COSMOS2020. Figure 6 shows the 5.6 m flux distributions of the COSMOS2020 undetected sources vs. the matched sources. Unsurprisingly, the unmatched sources are generally fainter than the overall MIRI source population. Indeed, visual examination of the 108 unmatched sources showed only 31 are unambiguous MIRI detections. Removing these sources which are clean MIRI detections despite being unmatched in COSMOS2020 leaves us with 77 likely fake sources ( of the total). Many of these apparently fake sources are driven by single bright pixels that were missed in the standard pipeline reduction, others appear to be residuals of the striping pattern seen in Figure 2. In Figure 6 we show separately the flux distributions of the unmatched sources judged to be real vs. those that are potentially fake. Note that the higher Section 4.4 explores in more detail the nature of the 31 reliable MIRI sources that are unmatched in COSMOS2020.
@@ 4.1 Source completeness
For the injected fake sources, we need a model of their on-the-sky distribution. Simply adopting the MIRI PSF, is not appropriate since the vast majority of our sources are spatially resolved. To create a more realistic source model, we stack 500 isolated and visually compact sources to construct a source model stamp. Figure 7 shows this stamp which measures 16 pixels 16 pixels (1.81.8′′). This source model stamp has a mean (along x- and y-axis) full width at half-maximum (FWHM) of 0.357′′ which, as expected, is larger than the MIRI 5.6 m PSF of 0.207′′, as measured during JWST commissioning444For details, refer to JWST user documentation at: https://jwst-docs.stsci.edu/jwst-mid-infrared-instrument/miri-performance/miri-point-spread-functions..
Completeness estimation starts with a segmentation map produced by SExtractor where all pixels with zero value correspond to no source present. We additionally set all pixel values to zero if they were masked, as described in Section 3.2, as well as setting to zero all sources consisting of 25 or fewer contiguous pixels. This second step was done in order to provide more sites suitable for injection, as our procedure avoids injecting in sites flagged by the segmentation map as having sources, though we deem this unnecessary for smaller sources. We then proceeded to inject fake sources across a range of flux densities ranging between 0.01 Jy and 0.25 Jy in 25 total increments. Sources are injected only on valid pixels not removed in the above steps. At each flux level, we injected 500 sources in the image. The sources are injected by using the source model scaled to the desired flux density. We then processed the newly-created images with the same SExtractor procedure used for the science images themselves. The ratio of recovered to injected sources at each flux density level constitutes the completeness at that level. This procedure was done 10 times per flux level per image, totaling 80 runs. Figure 8 shows the measured completeness curves for each run as well as the adopted completeness curve, which is the median between the 80 runs. As expected we find that the completeness is essentially 100% above our estimated 5 image depth and drops rapidly below that. The 50% completeness is at 0.07 Jy.
Right Our integral number counts for our eight fields (orange symbols), with the Poisson and field-to-field variation uncertainties added in quadrature. We overlay for comparison both the published CEERS counts (Yang et al., 2023, blue crosses, include only Poisson error) and the SMILES counts (Stone et al., 2024, green circles, include only Poisson error), as well as model predictions MIRI F560W counts from Cowley et al. (2018, black curves) where the dot-dash curve is their fiducial model and the upper (dashed) one is the same but with evolving supernovae feedback (“EvolFB”). We also overlay the limit as a vertical dashed line.
@@ 4.2 Number counts
We calculate the MIRI 5.6 m integral number counts by first counting the number of objects in each flux bin as . The source completeness for each flux bin is and is as shown in Figure 8. The integral counts are given by:
@@
We calculate the effective area, , as the inverse of the Boolean mask used for source detection (Section 3.2) times the single pixel area. For all 8 fields combined, this adds up to 18.4 sq.arcmin. Note that our coverage is fairly uniform within these masks therefore we assume the effective area does not change with flux level. The calculated integral counts are given in Table 3 along with their Poisson uncertainties. To calculate the raw counts, i.e. the ones not corrected for source comleteness, we simply adopt in all flux bins in Equation 1. When computing our number counts, we use less stringent quality cut on our catalog, counting all sources with an SNR . In Figures 9 we overlay the limit of 0.1 Jy on our counts figures as reference.
@@
Figure 9 left shows both our combined MIRI 5.6 m integral number counts, with Poisson errors, and the counts estimated for each of our 8 fields individually. It is obvious that the field-to-field variation is significantly greater than the Poisson uncertainties on the combined counts. The shape of the counts is fairly consistent with nearly all fields showing a knee at around 6-8 Jy.
@@
Figure 9 right shows the combined integral MIRI 5.6 m number counts where the uncertainties on the points represent the total uncertainties. The later are estimated by adding in quadrature the Poisson uncertainties with a field-to-field variation uncertainty which is taken to be half of the total spread between the individual field counts at each flux level. The combined counts and their Poisson and total uncertainties are given in Table 3.
@@
For comparison, in Figure 9 we overlay the CEERS 5.6 m counts from (Yang et al., 2023). The data for the CEERS number counts come from four nearby fields (separated by 15′) with a total area of sq.arcmin. The median depth for the CEERS images at 5.6 m is 0.138 Jy (Yang et al., 2023). We note that our counts tend to be higher than the CEERS ones, but this disagreement is only at 1 level for most flux bins considering our total uncertainties that account for field-to-field variation. Note that the uncertainties plotted on the CEERS counts are purely Poisson and do not include field-to-field uncertainties. The gap between our counts and their only becomes significant at the lowest flux bins, below 0.3 Jy; the widest part of the gap is . We also overlay the just released MIRI 5.6 m counts from the SMILES team (Stone et al., 2024). They are in good agreement with our counts at the bright end, but closer to CEERS at the faint end, though again within our 1 uncertainties. We note that both CEERS and SMILES use combined F560W+F770W images for source detection, unlike this analysis which is done entirely with MIRI 5.6 m images. To test the potential effect of differences in photometry, we ran our SExtractor setup on the public background-subtracted CEERS images and compared with their photometric catalog. We find excellent agreement, with the largest discrepancies per flux bin (e.g. for 1 Jy fluxes) being that we find roughly 10% more sources. This is consistent with our estimated fake source fraction in Section 3.4. Such fake sources are less likely in the CEERS photometric catalog due to their including the F770W as well for source detection. This accounts for some of the offset we see at fainter fluxes, but 10% likely fake source fraction is far from sufficient to explain the gap which is 2 at its maximum. Our Figure 9 left suggests this is primarily due to field-to-field variation instead.
@@
In Figure 9 right we also overlay the model counts from Cowley et al. (2018). Their predictions are based on the GALFORM semi-analytic model for galaxy formation within a CDM framework (Lacey et al., 2016). The predictions based on the core galaxy formation model from Lacey et al. (2016) are referred to as the ”Baseline model”. The model that includes supernovae feedback whose strength evolves with redshift is called the “EvolFB model”555We obtained the Cowley et al. (2018) model data from http://icc.dur.ac.uk/data/.. The GALFORM model assumes the Maraston (2005) stellar population models and the dust radiative transfer comes from the GRASIL prescription of Silva et al. (1998). Our counts, in the brighter regime, are reasonably consistent with these model predictions with a slight preference for their evolving feedback model, although the difference between their models is well within our uncertainaties (including the field-to-field variation). In agreement with the CEERS and SMILES results, we find that the Cowley et al. (2018) models overestimate the counts for sources below 3 Jy.
@@
4.3 The MIRI 5.6 m sources in the COSMOS2020 catalog
As already discussed in Section 3.3, we cross-matched the sources within our two COSMOS fields with the COSMOS2020 catalog (Weaver et al., 2022a, b). The overlap between our MIRI detections, the COSMOS2020 sources and their subset with IRAC channel 3 detections is illustrated in Figure 10 in the case of the COSMOS1 field. Note that with these new data more than half of the COSMOS2020 sources now have 5.6 m detections. The number of sources with MIRI 5.6 m detection in this field is more than 4 the sources with previous IRAC channel 3 (5.8 m) detections. This is illustrative of the fact that until JWST, it was difficult to link the optical and mid-IR source populations. Within our two COSMOS MIRI fields, there are 690 MIRI sources, of which 582 (84%) are matched to COSMOS2020 catalog sources, using a 1 matching radius (see Section 3.3).

### v2

@@ 2.5 Photometric zero-point
Pre-launch estimates of the MIRI imager’s performance already predicted that its point spread function (PSF) has a ‘cruciform’ artifact that draws luminosity from the core into the wings of the Gaussian. In-flight measurements have shown that this artifact is prominent at shorter wavelengths like ours (Rigby et al., 2023), but is negligible at (Gáspár et al., 2021). The extent of this effect was only noted once in-flight assessments had been made. Since our data were obtained very early after the start of science operations, the FITS header keyword PHOTUJA2 provided for converting from the native units to physical photometry units was incorrect. The fiducial keyword translated to a photometric zero-point of 26.121 mag which resulted in us seeing a significant systematic offset in the photometry measured by MIRI F560W as compared to the Spitzer IRAC Ch3 data available for our fields.
This issue was also noted by the CEERS team and led them to calculate a photometric zero-point for the images of 25.701 in AB magnitudes (Papovich et al., 2023). However, the CEERS images have a pixel scale of 0.09 (Papovich et al., 2023) whereas ours is 0.11. We re-scale the CEERS calculated zero-point as in . This is the value we adopt in this paper which leads to much better agreement between the MIRI and IRAC channel 3 photometry for cross-matched sources (see Section 3.3).
@@ 3.1 SExtractor photometry
We used Source Extractor (Bertin, E. & Arnouts, S., 1996) version 2.28.0 for source detection and photometry. Table 2 lists our SExtractor parameter settings. The SEEINGFWHM keyword adopted here corresponds to the FWHM of the in-flight measured MIRI point spread function (PSF) (see section 4.1 for details). As our fiducial flux measurements, we adopt values produced with Kron aperture photometry, labeled by the FLUX_AUTO keyword by SExtractor. The photometric uncertainties are natively provided by SE and take into account the variance of the pixel values inside the aperture, the background value, and the effective detector gain.
@@ 3.2 Masking
We further clean the raw SExtractor catalogs by masking out bright stars as well as any particularly noisy parts of the images. For each image, we created masks that include two components. One excludes obviously diffracted sources. The other avoids image edges555The edges are masked since, besides the issue of sources falling partly outside the image area, we have higher backgrounds there (see Figure 1) largely due to the lower exposure times on the edges since the primary MRS observations involved a small dither, see Young et al. (2023, in prep). where we assume an edge width of 5 pixels. The masks also include the blocked regions of the Lyot coronograph) and the residual striping artifacts – most of these were removed in the background subtraction step but some residual striping remained near the bottom of the images, which we define manually (see right-hand panel in Figure 2). Figure 3 shows as an example the mask for the FLS1 field.
We also apply a quality cut on our catalogs, only keeping sources with a signal-to-noise ratio greater than 5. The combination of the masking and the quality cut reduces the raw SExtractor output of 4,089 sources to a total of 2,854 “reliable” sources (“reliable” meaning falling outside our defined masks and having SNR) among all our eight fields. The masking improves the reliability of our source population, since within the masked areas we find a significantly larger fraction of fake sources, relative to the rest of the images as discussed in Section 3.4, due to “sources” being picked up by SExtractor on top of strong diffraction spikes or stripes. Per field, after masking we detect the most sources in field FLS3 (409 sources) and the least in field FLS5 (310 sources).
@@ 3.3 Astrometric & photometric verification
In order to perform astrometric and photometric verifications, we used the SExtractor detected sources in our two COSMOS fields and cross-matched them with the COSMOS2020 CLASSIC catalog (Weaver et al., 2022a). The COSMOS2020 source detection is based on a weighted combined image (Weaver et al., 2022a). We used a matching radius and found a total of 582 COSMOS2020 cross-matches, 175 of which also have IRAC channel 3 detections. We choose a radius as reasonable in a cross-match with the COSMOS2020 ground-based optical to near-IR data. As we see below, the vast majority of the matched sources have positional separations within so increasing the matching radius would not alter the results significantly while increasing the likelihood of spurious matches. Overall, our MIRI data increase the fraction of COSMOS2020 with 5.6/ data by roughly 3 times. Note that our MIRI fields are within the deep part of the COSMOS2020 catalog and not near the edges and the COSMOS2020 positions rely on Gaia astrometric solutions (Weaver et al., 2022a). Figure 4 shows the comparison between the MIRI and COSMOS2020 source positions of the cross-matched sources. The offsets we find are small and well within the MIRI PSF at this wavelength. Therefore we do not explicitly apply astrometric corrections.
Figure 5 shows the comparison between the MIRI and the available IRAC channel 3 fluxes. As described in section 3.1, we use Kron aperture photometry for the fluxes. For the IRAC fluxes, we used the aperture corrected aperture photometry from Weaver et al. (2022b). We overlay the 1:1 line. Overall, we find good agreement, although we have a median offset of 0.22 magnitudes. We compare this to the results presented in Figure 8 of Yang et al. (2023), who have significantly deeper IRAC data and therefore are able to compute the median offset in magnitude bins down to 25th magnitude. Their median offset at their IRAC 5 bin of 22.5 magnitude is roughly 0.10.1, and gets much larger at fainter magnitudes and lower SNR. Their offset is near zero at magnitudes brighter than 21. Most of our sources with IRAC SNR3 are in the 21-22.5 magnitude range. Here we are consistent with Yang et al. (2023) in that we find a positive offset (i.e. generally the MIRI magnitudes are fainter) and our median offset is within 1 of theirs for the 22.5 magnitude bin.
@@ 3.4 Estimating fraction of fake sources
The cross-match with COSMOS2020 also gives us an upper limit on the fraction of the SExtractor detected sources that might be fake. Such sources are expected due to some of the residual artifacts in the images, that may have been missed in the masks described in Section 3.2. Overall we found COSMOS2020 counterparts for 582 of the 690 MIRI sources in our two COSMOS fields (84 %). The 108 unmatched sources (), represent an upper limit on the fake source fraction since it is expected that we have some red MIRI sources that are not detected in COSMOS2020. Figure 7 shows the flux distributions of the sources matched to COSMOS2020, compared to those of unmatched sources. Unsurprisingly, the unmatched sources are generally fainter than the overall MIRI source population. Indeed, visual examination of the 108 unmatched sources showed only 31 are unambiguous MIRI detections. Removing these sources which are clean MIRI detections despite being unmatched in COSMOS2020 leaves us with 77 likely fake sources ( of the total). Some of these apparently fake sources are driven by single bright pixels that were missed in the standard pipeline reduction, others appear to be residuals of the striping pattern seen in Figure 2 or other image artifacts. Figure 6 shows examples of sources we flagged as fake. In Figure 7 we show separately the flux distributions of the unmatched sources judged to be real vs. those that are potentially fake. Note that Section 4.4 explores in more detail the nature of the 31 reliable MIRI sources that are unmatched in COSMOS2020.
@@ 4.1 Source completeness
For the injected fake sources, we need a model of their on-the-sky distribution. Simply adopting the MIRI PSF, is not appropriate since the vast majority of our sources are spatially resolved. To create a more realistic source model, we stack 500 isolated and visually compact sources to construct a source model. Figure 8 shows this model which measures 16 pixels 16 pixels (). This source model has a mean (along x- and y-axis) full width at half-maximum (FWHM) of which, as expected, is larger than the MIRI m PSF of , as measured during JWST commissioning666For details, refer to JWST user documentation at: https://jwst-docs.stsci.edu/jwst-mid-infrared-instrument/miri-performance/miri-point-spread-functions..
Completeness estimation starts with a segmentation map produced by SExtractor where all pixels with zero value correspond to no source present. We additionally set all pixel values to zero if they were masked, as described in Section 3.2, as well as setting to zero all sources consisting of 25 or fewer contiguous pixels. This second step was done in order to provide more sites suitable for injection, as our procedure avoids injecting in sites flagged by the segmentation map as having sources, though we deem this unnecessary for smaller sources. We then proceeded to inject fake sources across a range of flux densities ranging between and in 25 total increments. Sources are injected only on valid pixels not removed in the above steps. At each flux level, we injected 500 sources in the image. The sources are injected by using the source model scaled to the desired flux density. We then processed the newly-created images with the same SExtractor procedure used for the science images themselves. The ratio of recovered to injected sources at each flux density level constitutes the completeness at that level. This procedure was done 10 times per flux level per image, totaling 80 runs. Figure 9 shows the measured completeness curves for each run as well as the adopted completeness curve, which is the median between the 80 runs. As expected we find that the completeness is essentially 100% above our estimated 5 image depth and drops rapidly below that. The 50% completeness is at .
@@ 4.2 Number counts
We calculate the MIRI integral number counts by first counting the number of objects in each flux bin as . The source completeness for each flux bin is and is as shown in Figure 9. The integral counts are given by:
@@
We calculate the effective area, , as the inverse of the Boolean mask used for source detection (Section 3.2) times the single pixel area. For all eight fields combined, this adds up to 18.4 square arcminutes. Note that our coverage is fairly uniform within these masks therefore we assume the effective area does not change with flux level. The calculated integral counts are given in Table 3 along with their Poisson uncertainties. To calculate the raw counts, i.e. the ones not corrected for source completeness, we simply adopt in all flux bins in Equation 1. When computing our number counts, we use less stringent quality cut on our catalog, counting all sources with an SNR . In Figures 10 we overlay the limit of on our counts figures as a reference point.
@@
The left panel of Figure 10 shows our combined MIRI integral number counts, with Poisson errors, along with the counts estimated for each of our eight fields individually. It is obvious that the field-to-field variation is significantly greater than the Poisson uncertainties on the combined counts. The shape of the counts is fairly consistent with nearly all fields showing a knee at around 6-.
@@
The right panel of Figure 10 shows the combined integral MIRI number counts where the uncertainties on the points represent the total uncertainties. The latter are estimated by adding in quadrature the Poisson uncertainties with a field-to-field variation uncertainty which is taken to be half of the total spread between the individual field counts at each flux level. The combined counts and their Poisson and total uncertainties are given in Table 3.
@@
For comparison, in Figure 10 we overlay the CEERS counts from (Yang et al., 2023). The data for the CEERS number counts come from four nearby fields (separated by ) with a total area of square arcminutes. The median depth for the CEERS images at is (Yang et al., 2023). We note that our counts tend to be higher than the CEERS ones. This disagreement is only at 1 level for most flux bins considering our total uncertainties that account for field-to-field variation. The offset between our counts and the CEERS ones suggests the CEERS fields might be slightly underdense. Our widely separated fields allow us both an estimate of the average counts as well as a sense of the field-to-field variation. For example, in the left panel of Figure 10, some of our fields do show number counts consistent with those in the CEERS fields. Note that the uncertainties plotted on the CEERS counts are purely Poisson and do not include field-to-field uncertainties. The gap between our counts and theirs only becomes significant at the lowest flux bins, below ; the widest part of the gap is . We also overlay the just released MIRI counts from the SMILES team (Stone et al., 2024). They are in good agreement with our counts at the bright end, but closer to CEERS at the faint end, though again within our 1 uncertainties. We note that both CEERS and SMILES use combined F560W and F770W images for source detection, unlike this analysis which is done entirely with MIRI images. To test the potential effect of differences in photometry, we ran our SExtractor setup on the public background-subtracted CEERS images and compared with their photometric catalog. We find excellent agreement, with the largest discrepancies per flux bin (e.g. for fluxes) being that we find roughly 10% more sources. This is consistent with our estimated fake source fraction in Section 3.4. Such fake sources are less likely in the CEERS photometric catalog due to their including the F770W data as well for source detection. This accounts for some of the offset we see at fainter fluxes, but 10% likely fake source fraction is far from sufficient to explain the gap which is 2 at its maximum. Our Figure 10 left suggests this is primarily due to field-to-field variation instead.
@@
In the right panel of Figure 10 we also overlay the model counts from Cowley et al. (2018). Their predictions are based on the GALFORM semi-analytic model for galaxy formation within a CDM framework (Lacey et al., 2016). The predictions based on the core galaxy formation model from Lacey et al. (2016) are referred to as the ”Baseline model”. The model that includes supernovae feedback whose strength evolves with redshift is called the “EvolFB model”777We obtained the Cowley et al. (2018) model data from http://icc.dur.ac.uk/data/.. The GALFORM model assumes the Maraston (2005) stellar population models and the dust radiative transfer comes from the GRASIL prescription of Silva et al. (1998). Our counts, in the brighter regime, are reasonably consistent with these model predictions with a slight preference for their evolving feedback model, although the difference between their models is well within our uncertainaties (including the field-to-field variation). In agreement with the CEERS and SMILES results, we find that the Cowley et al. (2018) models overestimate the counts for sources below .
@@
4.3 The MIRI sources in the COSMOS2020 catalog
As already discussed in Section 3.3, we cross-matched the sources within our two COSMOS fields with the COSMOS2020 catalog (Weaver et al., 2022a, b). The overlap between our MIRI detections, the COSMOS2020 sources and their subset with IRAC channel 3 detections is illustrated in Figure 11 in the case of the COSMOS1 field. Note that with these new data more than half of the COSMOS2020 sources now have detections. The number of sources with MIRI detection in this field is more than 4 the sources with previous IRAC channel 3 () detections. This is illustrative of the fact that until JWST, it was difficult to link the optical and mid-IR source populations. Within our two COSMOS MIRI fields, there are 690 MIRI sources, of which 582 (84%) are matched to COSMOS2020 catalog sources, using a 1 matching radius (see Section 3.3).